# Tensor-product interactions in Markov-switching models

### v2

@@ Abstract
Markov-switching models are a powerful tool for modelling time series data that are driven by underlying latent states. As such, they are widely used in behavioural ecology, where discrete states can serve as proxies for behavioural modes and enable inference on latent behaviour driving e.g. observed movement. To understand drivers of behavioural changes, it is common to link model parameters to covariates. Over the last decade, nonparametric approaches have gained traction in this context, particulary to avoid unrealistic parametric assumptions. Nonetheless, existing methods are largely limited to univariate smooth functions of covariates, based on penalised splines, while real processes are typically complex requiring consideration of interaction effects. We address this gap by incorporating tensor-product interactions into Markov-switching models, enabling flexible modelling of multidimensional effects in a computationally efficient manner. Based on the extended Fellner-Schall method, we develop an efficient automatic smoothness selection procedure that is robust and scales well with the number of smooth functions in the model. The method builds on a random effects view of the spline coefficients and ultimately yields a recursive penalised likelihood procedure. As special cases, this general framework accommodates bivariate smoothing, function-valued random effects, and space-time interactions. We demonstrate its practical utility through three ecological case studies of an African elephant, common fruitflies, and Arctic muskoxen. The methodology is fully implemented in the LaMa R package, providing applied ecologists with an accessible and flexible tool for semiparametric inference in hidden-state models. The approach has the potential to drastically improve the level of detail in inference, allowing to fit HMMs with hundreds of parameters, 10-20 (potentially bivariate) smooths to thousands of observations.
@@ 3 Tensor-product interactions and associated penalties
$-\lambda\bm{\beta}^{\intercal}\bm{S}\bm{\beta}$
to the log-likelihood function, where $\bm{S}$ is a fixed penalty matrix that depends on the basis used and our operationalisation of smoothness. For the two-dimensional basis of interest here, we could, for example, choose a thin-plate regression spline basis, which would yield the fixed basis functions $B_{k}$ and associated penalty matrix $\bm{S}$. While incorporating a two-dimensional function, such a predictor still yields a “simple" smooth with a single smoothing parameter and could be estimated using the tools developed by Michelot (2022) or Koslik (2024).
@@ 4 Smoothness selection via the extended Fellner-Schall method
As discussed in the previous section, a tensor-product smooth can be represented either as a single smooth with two penalty matrices (hence also two penalty parameters) or as a combination of two “simple" smooths along with a pure interaction term. In general, a Markov-switching model may include several simple smooths as well as several tensor-product interactions — potentially one for each state or each off-diagonal entry of the t.p.m. For notational simplicity, it is thus helpful to represent the zoo of distinct penalties on various subvectors of the coefficient vector $\bm{\theta}$ by a single penalty on the entire coefficient vector. Consequently, we define the full-model penalty matrix
@@
[…]
@@
$\ell_{r}(\bm{\lambda})=\ell(\hat{\bm{\theta}}_{\lambda})+\hat{\bm{\theta}}_{\lambda}^{\intercal}\bm{S}_{\lambda}\hat{\bm{\theta}}_{\lambda}/2+\log|\bm{S}_{\lambda}|_{+}/2-\log|\bm{H}_{\lambda}+\bm{S}_{\lambda}|/2+\text{const.},$. , (6)
@@
Hence, a further approximation is needed that allows for a good tradeoff between estimation accuracy and computational efficiency. A good candidate is given by the so-called extended Fellner-Schall method (Fellner, 1986; Schall, 1991; Wood and Fasiolo, 2017), which proceeds as follows. Partially differentiating (6) w.r.t. $\lambda_{j}$ — treating $\hat{\bm{\theta}}_{\lambda}$ as a constant — yields
$\frac{\partial\ell_{r}}{\partial\lambda_{j}}=\hat{\bm{\theta}}_{\lambda}^{\intercal}\mathcal{S}_{j}\hat{\bm{\theta}}_{\lambda}/2+\text{tr}(\bm{S}_{\lambda}^{-}\mathcal{S}_{j})/2-\text{tr}(\bm{J}_{\lambda}^{-1}\mathcal{S}_{j})/2-\text{tr}(\bm{J}_{\lambda}^{-1}\partial\bm{H}_{\lambda}/\partial\lambda_{j})/2,$ (7)
where $\bm{J}_{\lambda}=\bm{H}_{\lambda}+\bm{S}_{\lambda}$. The term directly involving the log-likelihood vanishes, since $\partial\ell/\partial\bm{\theta}|_{\hat{\bm{\theta}}_{\lambda}}=\bm{0}$ by definition of $\hat{\bm{\theta}}_{\lambda}$. The most problematic part of the above equation is the last term, as it involves the explicit dependence of the Hessian of the negative log-likelihood on the smoothing parameter. However, Breslow and Clayton (1993); Gu (1992); Wood and Fasiolo (2017) neglect the dependence of $\bm{H}_{\lambda}$ on $\bm{\lambda}$ justified by it vanishing asymptotically. Having made this approximation, (7) (without its last term) can now be used to construct the estimating equation
@@ 5.1 African elephant — year-round diurnal variation
To these data, we fitted 2-state hidden Markov models (HMMs), assuming state-dependent gamma distributions for the step lengths and von Mises distributions for the turning angles, with conditional independence between the two observed variables given the underlying state. While a 2-state model may be an overly simplistic representation of the elephant’s behaviour, the relatively coarse temporal resolution does not allow for more detailed inference regarding its behavioural process. An initial fit with a homogeneous Markov chain as the state-process model suggested that the first state is characterised by slow, undirected movement, whereas the second state involved larger steps and more directed movement. Accordingly, we interpret the two states as “encamped" and “exploratory" behaviour (cf. Morales et al., 2004).
@@
s(tday, bs = ’cc’, k = 12) + s(julian, bs = ’cc’, k = 12) +
ti(tday, julian, bs = ’cc’, k = 12).
@@
In total, the model with the tensor-product interaction comprises 248 coefficients (the cyclic bases results in $12010+10+10^{2}=$ 120 coefficients for each linear predictor) and eight smoothness parameters that need to be estimated — four for each off-diagonal entry of the t.p.m. These smoothness parameters were initialised with $10^{4}$ for the main effects and $10^{5}$ for the interaction. Model estimation then took about seven minutes on an Apple M2 chip with 16 GB of memory. The extended Fellner-Schall method required a total of 22 penalised fits for convergence of the outer optimisation, i.e. until the largest component of the approximate outer gradient fell below $10^{-4}$. Estimation of the other candidate models was substantially faster, taking under one minute each, so we do not report the exact times here.
@@ 5.2 Drophila melanogaster — function-valued random effects
[…]
@@
By expressing the function-valued random effect using the ANOVA decomposition of a tensor-product interaction, separate variance parameters need to be estimated for the random intercepts $\alpha_{a,c}^{(ij)}$ and the “random-effect dimension" of $f_{c,a}^{(ij)}(\text{tday}_{t})$. Indeed, this is desirable for biological understanding: The first variance parameter quantifies the variation in the overall activity level, also called behavioural type, while the second variance parameter quantifies heterogeneity in the behavioural response to the time of day, also called behavioural plasticity (Hertel et al., 2020).

### v3

@@ Abstract
Markov-switching models are a powerful tool for modelling time series data that are driven by underlying latent states. As such, they are widely used in behavioural ecology, where discrete states serve as proxies for behavioural modes and enable inference on latent behaviour driving e.g. observed movement. To understand drivers of behavioural changes, it is common to link model parameters to covariates, with nonparametric approaches having gained traction in this context to avoid unrealistic parametric assumptions. Existing methods are largely limited to univariate smooth functions of covariates while real processes are typically complex, requiring interaction effects. We address this gap by incorporating tensor-product interactions into Markov-switching models, enabling flexible modelling of multidimensional effects in a computationally efficient manner. Based on the extended Fellner-Schall method, we develop an automatic smoothness selection procedure that is robust and scales well with the number of smooth functions. The method builds on a random effects view of the spline coefficients and yields a recursive penalised likelihood procedure. As special cases, this general framework accommodates bivariate smoothing, function-valued random effects, and space-time interactions. We demonstrate its practical utility through ecological case studies of an African elephant, common fruitflies, and Arctic muskoxen. The methodology is implemented in the LaMa R package, providing applied ecologists with an accessible and flexible tool to fit models with hundreds of parameters and 10-20 (potentially bivariate) smooths.
@@ 3 Tensor-product interactions and associated penalties
$\lambda\bm{\beta}^{\intercal}\bm{S}\bm{\beta}$
to the log-likelihood function, where $\bm{S}$ is a fixed penalty matrix that depends on the basis used and our operationalisation of smoothness. For the two-dimensional basis of interest here, we could, for example, choose a thin-plate regression spline basis, which would yield the fixed basis functions $B_{k}$ and associated penalty matrix $\bm{S}$. While incorporating a two-dimensional function, such a predictor still yields a “simple” smooth with a single smoothing parameter and could be estimated using the tools developed by Michelot (2022) or Koslik (2024).
@@ 4 Smoothness selection via the extended Fellner-Schall method
As discussed in the previous section, a tensor-product smooth can be represented either as a single smooth with two penalty matrices (hence also two penalty parameters) or as a combination of two “simple” smooths along with a pure interaction term. In general, a Markov-switching model may include several simple smooths as well as several tensor-product interactions — potentially one for each state or each off-diagonal entry of the t.p.m. For notational simplicity, it is thus helpful to represent the zoo of distinct penalties on various subvectors of the coefficient vector $\bm{\theta}$ by a single penalty on the entire coefficient vector. Consequently, we define the full-model penalty matrix
@@
$\mathcal{L}_{r}(\bm{\lambda})=p_{\lambda}(\bm{x})=\int_{\mathbb{R}^{d}}p(\bm{x}\mid\bm{\theta})\>p_{\lambda}(\bm{\theta})\;d\bm{\theta},$ (5)
@@
$\ell_{r}(\bm{\lambda})=\ell(\hat{\bm{\theta}}_{\lambda})-\hat{\bm{\theta}}_{\lambda}^{\intercal}\bm{S}_{\lambda}\hat{\bm{\theta}}_{\lambda}/2+\log|\bm{S}_{\lambda}|_{+}/2-\log|\bm{H}_{\lambda}+\bm{S}_{\lambda}|/2+\text{const.},$ (6)
@@
Hence, a further approximation is needed that allows for a good tradeoff between estimation accuracy and computational efficiency. A good candidate is given by the so-called extended Fellner-Schall method (Fellner, 1986; Schall, 1991; Wood and Fasiolo, 2017), which proceeds as follows. Partially differentiating (6) w.r.t. $\lambda_{j}$ yields
$\frac{\partial\ell_{r}}{\partial\lambda_{j}}=-\hat{\bm{\theta}}_{\lambda}^{\intercal}\mathcal{S}_{j}\hat{\bm{\theta}}_{\lambda}/2+\text{tr}(\bm{S}_{\lambda}^{-}\mathcal{S}_{j})/2-\text{tr}(\bm{J}_{\lambda}^{-1}\mathcal{S}_{j})/2-\text{tr}(\bm{J}_{\lambda}^{-1}\partial\bm{H}_{\lambda}/\partial\lambda_{j})/2,$ (7)
where $\bm{J}_{\lambda}=\bm{H}_{\lambda}+\bm{S}_{\lambda}$. The term directly involving the log-likelihood vanishes, since $\partial\ell_{p}/\partial\bm{\theta}|_{\hat{\bm{\theta}}_{\lambda}}=\bm{0}$ by definition of $\hat{\bm{\theta}}_{\lambda}$. The most problematic part of the above equation is the last term, as it involves the explicit dependence of the Hessian of the negative log-likelihood on the smoothing parameter. However, Breslow and Clayton (1993); Gu (1992); Wood and Fasiolo (2017) neglect the dependence of $\bm{H}_{\lambda}$ on $\bm{\lambda}$ justified by it vanishing asymptotically. Having made this approximation, (7) (without its last term) can now be used to construct the estimating equation
@@ 5.1 African elephant — year-round diurnal variation
To these data, we fitted 2-state hidden Markov models (HMMs), assuming state-dependent gamma distributions for the step lengths and von Mises distributions for the turning angles, with conditional independence between the two observed variables given the underlying state. While a 2-state model may be an overly simplistic representation of the elephant’s behaviour, the relatively coarse temporal resolution does not allow for more detailed inference regarding its behavioural process. An initial fit with a homogeneous Markov chain as the state-process model suggested that the first state is characterised by slow, undirected movement, whereas the second state involved larger steps and more directed movement. Accordingly, we interpret the two states as “encamped” and “exploratory” behaviour (cf. Morales et al., 2004).
@@
s(tday, bs = "cc", k = 12) + s(julian, bs = "cc", k = 12) +
ti(tday, julian, bs = "cc", k = 12).
@@
In total, the model with the tensor-product interaction comprises 248 coefficients (the cyclic bases results in $10+10+10^{2}=120$ coefficients for each linear predictor) and eight smoothness parameters that need to be estimated — four for each off-diagonal entry of the t.p.m. These smoothness parameters were initialised with $10^{4}$ for the main effects and $10^{5}$ for the interaction. Model estimation then took about seven minutes on an Apple M2 chip with 16 GB of memory. The extended Fellner-Schall method required a total of 22 penalised fits for convergence of the outer optimisation, i.e. until the largest component of the approximate outer gradient fell below $10^{-4}$. Estimation of the other candidate models was substantially faster, taking under one minute each, so we do not report the exact times here.
@@ 5.2 Drophila melanogaster — function-valued random effects
For this case study, we compare the wild type to one of the 13 originally studied modified genotypes, namely $\text{Hsp83}^{\text{e6A}}$ / $\text{Hsp83}^{\text{08445}}$. The two groups contain 35 and 34 individuals, respectively. In total, the data comprises 13,104 observations of the wild-type and 13,056 observations of the modified geneotype.
@@
By expressing the function-valued random effect using the ANOVA decomposition of a tensor-product interaction, separate variance parameters need to be estimated for the random intercepts $\alpha_{a,c}^{(ij)}$ and the “random-effect dimension” of $f_{c,a}^{(ij)}(\text{tday}_{t})$. Indeed, this is desirable for biological understanding: The first variance parameter quantifies the variation in the overall activity level, also called behavioural type, while the second variance parameter quantifies heterogeneity in the behavioural response to the time of day, also called behavioural plasticity (Hertel et al., 2020).